OpenStudy (anonymous):

solve 5/8x+1/16x=5/16+x

OpenStudy (anonymous):

First find the LCM Which is 16 this case Multiply through by 16 16(5/8)x + 16(1/16)x = 16(5/16) + 16(x) Simplify get 10x + x = 5 + 16x Subtract 16x from both sides of the equation 10x + x -16x = 5 + 16x - 16x Combine like terms -5x = 5 Divide both sides by -5 (-5x) / (-5) = 5 / (-5) x = -1